# Model Provenance Testing for Large Language Models

### v1

@@ 3.2 Model Provenance Tester
$\mathcal{M}\leftarrow\{\mu_{1},\ldots,\mu_{s}\}\cup\{\mu^{\prime}_{1},\ldots,\mu^{\prime}_{m}\}$triangleright Set of all sims
$\mu_{max}\leftarrow\max(\mathcal{M})$triangleright Find highest sim
@@
     $p_{i}\leftarrow\text{ZTest}(\mu_{max},\mu^{\prime},T)$triangleright Compare against other sims
@@
for $k\leftarrow 1$ to $s+m-1$ do
     $\alpha_{k}\leftarrow\alpha/(s+m-k)$triangleright Holm-Bonferroni adjustment
@@
end forreturn $(\textsc{True},\operatorname*{arg\,max}_{i\in[s]}\mu_{i})$triangleright Return parent
Extended Model Provenance Test.
The above provenance problem described thus far assumes we have one candidate parent $f$ that we want to test our child model $g$ against. We now extend it to the general case when there is a set of parents, which we refer to as the extended model provenance problem (with unspecified parent). In this problem, given only query access to models, the goal is to determine whether a model $g$ is derived from some model from the set ${f_{1},\ldots,f_{s}}$ of candidate parent models.
While running the basic tester $s$ times (once for each provenance pair $(f_{i},g)$) would solve the extended parent problem, this approach besides requiring more effort, also would require additional correction for multiple testing to maintain the same level of confidence. The probability of false positives would grow with the number of candidate parents $s$ unless appropriate adjustments (such as Holm-Bonferroni correction) are made to the significance level. We thus consider improved tester given in Algorithm 2. Our tester avoids this issue by conducting a single set of hypothesis tests after identifying the most similar candidate. It works as follows. First it finds the most similar model to the given model $g$ among all the control models $C$ and candidate parents $F$. If that model is a control model, the algorithm terminates with False. Otherwise, it goes on to test whether the FWER of this model is overall below $\alpha$, the desired significance level. The test for the latter is the same as in Algorithm 1, except now all the alternate hypotheses (including control and candidates) are in the family tested against. When the algorithm return True, it has the guarantee that the most similar model is one of the candidate models and that the total significance level across all hypotheses meets the threshold $\alpha$.
@@ 3.3 Understanding Sources of Error
Our provenance tester can make two types of errors: false positives, where it incorrectly identifies a provenance relationship between independently developed models, and false negatives, where it fails to detect an actual derivation relationship between models.
These errors arise in part from our statistical hypothesis tests. When performing such tests, we can control false positives by setting a stricter significance level $\alpha$ (requiring stronger evidence before claiming excess similarity), while we can reduce false negatives by increasing the sample size $T$ (more samples provide better power to detect true similarities).
Besides errors introduced by statistical tests, our approach of testing provenance through similarity detection may introduce additional errors. Since we claim provenance only when we detect higher-than-expected similarity between models, we need to examine what this approach implies for our error analysis. This approach relies on two key assumptions that can impact error rates:
Assumption 1: Derivation implies similarity.
We assume that when model $g$ is derived from $f$, they will exhibit above-average similarity in their outputs. This assumption leads to two potential types of errors:
  • False negatives occur when a derived model shows insufficient similarity to its parent. This can happen when a model customizer applies extensive modifications that significantly alter the model’s behavior. While resource constraints typically prevent such extreme modifications (as they would approach the cost of training from scratch), some legitimate derivation relationships may still go undetected.
  • False positives arise when independently developed models exhibit high similarity. This typically happens when models are trained on similar datasets or designed for similar specialized tasks - for instance, two independent medical diagnosis models may produce very similar outputs due to their shared domain constraints.
Assumption 2: Control models establish a valid baseline.
We assume our control models provide a reliable baseline for the similarity we should expect between unrelated models. Poor selection of control models can lead to two types of errors:
  • False positives occur when our control models are too dissimilar from the domain of $f$. For example, using general language models as controls for specialized code generation models sets an artificially low baseline, making normal domain-specific similarities appear significant.
  • False negatives happen when control models are themselves derived from $f$ or trained on very similar data. This establishes an artificially high baseline that masks genuine derivation relationships.
The overall error rates of our tester depend on the combination of errors from both our statistical hypothesis testing and the two core assumptions. While we can provide theoretical guarantees for controlling error rates in hypothesis testing through parameters $\alpha$ and $T$, we cannot derive analytical bounds for errors arising from the assumptions about derivation implying similarity or the validity of the control model baseline. These assumption-based error rates can only be evaluated empirically. However, our extensive experiments in Section 4 demonstrate that these assumptions hold well in practice across a wide range of models, suggesting that our approach of provenance testing to similarity detection is sound for real-world applications with non-adaptive adversaries.
@@ 3.4 Reducing Query Complexity
Most of LLMs available currently allow cheap (even free) API access, thus the monetary query cost of running our testers is insignificant. When this is not the case, for example, either when the cost of queries is high (e.g. one query to OpenAI model O1 can cost more than \$1 [32]), or the models have some rate restrictions, one can consider enhancements to our testers from Algorithms 12. Furthermore, there are use cases when query complexity can be reduced without any side effects, thus it makes sense from optimization perspective. Note that our proposed enhancements for query reduction are not meant to preserve the theoretical guarantees of classical hypothesis testing that our previous testers inherit, but they can be useful in setups where query costs are prohibitive.
@@
We can divide the queries used in the tester (see Algorithms 12) into two distinctive types: online queries made to the tested child model $g$, and offline queries made to the parent model $f$ (or models $f_{1},\ldots,f_{s}$) and to the control models $c_{1},\ldots,c_{m}$. We make this distinction for two reasons. First, often offline queries are much cheaper, as the potential parent models (and the control models as we will see in the Section 4) are well established, and available from multiple sources, thus they are usually cheaper or free. Second, in some use cases, we can reuse the offline queries to perform many provenance tests of different $g_{i}$. Thus further we analyze separately these two scenarios.
@@
Reducing Online Complexity. Since our tester is fundamentally based on statistical hypothesis testing, any reduction in query complexity must be compensated by increasing the statistical power of individual queries. Rather than querying model $g$ with $T$ random prompts, we can strategically select a smaller set of $T^{\prime}<T$ prompts that yield comparable statistical power for detecting model provenance [^2] It means in Algorithms 12, instead of random sampling $x_{1},\ldots,x_{T}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\Omega$, the goal is to find set $x_{1},\ldots,x_{T^{\prime}}$ from $x_{1},\ldots,x_{T}$ and $F,C$.. We achieve this through an informed sampling approach: instead of uniform sampling from $\Omega$, we employ rejection sampling with an entropy-based selection criterion. Specifically, to generate each prompt in $T^{\prime}$, we sample $k$ candidate prompts from $\Omega$ and select the one that maximizes the entropy of output tokens across all parent and control models. The selection criterion is dynamically weighted to favor prompts that have stronger discriminative power between similar models. While this approach introduces dependencies between the sampled prompts (so the theoretical guarantees of classical hypothesis testing used in Algorithm 1 and 2 do not carry over), our empirical results in Section 4.3 demonstrate its practical effectiveness. Full details about the approach are given in Appendix E.
Reducing Offline Complexity. In non-adversarial settings where multiple provenance tests are performed against the same parent model $f$, we can trivially reduce the offline complexity by reusing the same set of offline queries across all tests. A concrete example of such scenarios arises when issues are discovered in a pre-trained LLM, such as problematic training data or generation of harmful content. A recent example is the lawsuit against the pre-trained model LLama for using copyrighted data in its training set [7]. Since various teams and organizations may have fine-tuned their applications using this model, but precise provenance information is not readily available, there is a need to identify which models are derived from this problematic base model. In this case, the same set of offline queries to the base model and control models can be reused across all provenance tests.
@@
We further consider the case of reducing offline complexity in settings where offline queries cannot be reused. The current version of our provenance tester samples $T$ prompts for each parent/control model, then runs the hypothesis test to discover the most similar candidate to the tested model $g$ and shows it has significantly higher similarity. The key observation for reducing offline query complexity is that we may not need an equal number of queries to all parent/control models to identify the most similar one. If a particular parent model shows consistently higher similarity to $g$ compared to other models, we might be able to confirm it as the top candidate with fewer queries to the clearly dissimilar models. The challenge lies in determining when we have sufficient statistical evidence to conclude that one model is significantly more similar than the others, while maintaining our desired confidence levels.
This observation naturally leads us to formulate the problem as a Best Arm Identification (BAI) [5] problem in the Multi-Armed Bandit (MAB) setting. In this formulation, each parent or control model represents an “arm” of the bandit, and querying a model with a prompt corresponds to “pulling” that arm. The “reward” for each pull is the binary outcome indicating whether the model’s output matches that of the tested model $g$. The goal is to identify the arm (model) with the highest expected reward (similarity to $g$) while minimizing the total number of pulls (queries). So, we can leverage well-studied MAB algorithms that adaptively allocate queries, focusing more on promising candidates while quickly eliminating clearly dissimilar ones. The implementation of the tester based on BAI is detailed in Appendix F. Theoretical guarantees from the MAB literature could be applied to bound the number of queries needed to identify the correct parent model with high probability, but this is beyond our goals.
@@ 4 Evaluation
We evaluate our proposed provenance testing approach experimentally. Our evaluation aims to assess both the effectiveness of the approach and examine the validity of its core two assumptions. Specifically, we seek to answer the following research questions:
@@
Experimental Setup.
Models and Provenance Pairs.
We collect model candidates for all provenance pairs from the Hugging Face (HF) platform[21]. To avoid selection bias, we used download counts as our selection criterion, taking the most popular models subject only to hardware constraints on model size. To increase variety of candidates, we create two distinct benchmarks Bench-A and Bench-B, that differ in aspects such as model sizes, choice of pre-trained models, and ground-truth verification procedure. The full procedure of collection of models and constructions of benchmarks is described in Appendix A and their brief comparison is given in Table 1.
@@
model parameters 1B-4B $<$ 1B
compilation method manual (partially) automatic
@@
Framework.
Our default evaluation framework assumes extended model provenance testing (refer to Algorithm 2) as it allows to test for all parents at once. For sanity check, we also include an alternative framework, that considers the case where one parent is suspected (based on Algorithm 1), and this is discussed at the end of Section 4.1 and evaluated in Appendix D. We use the common value for significance parameter $\alpha=0.05$. Sampling of prompts is described in Appendix B.
Selection of control set.
In all of our provenance tests, we use the complete set of available pre-trained models from the benchmark as control models - 10 models for Bench-A and 57 for Bench-B. This selection was done to demonstrates that effective control sets can be constructed without careful manual curation or domain-specific analysis. Specifically, we make no effort to align control models with particular parent models’ domains or capabilities. We neither analyze the outputs of parent models $f$ nor attempt to match control models to specific use cases. Instead, we simply include all pre-trained models that rank among the most popular on the Hugging Face platform. This sampling approach, while simple, helps avoid introducing obvious selection bias while ensuring broad coverage of model types and capabilities. This straightforward selection strategy allows us to evaluate whether provenance testing can be effective even without carefully chosen control sets.
@@ 4.1 Accuracy of Model Provenance Tester
We evaluate the accuracy of the provenance tester by examining its performance on both Bench-A and Bench-B under different numbers of prompts. Figure 2 shows the precision and recall results from these experiments. The tester demonstrates similar performance patterns on both benchmarks, with slightly better results on Bench-A.
The precision is notably high (approximately $0.95$) when the tester uses up to $1,000$ prompts. Interestingly, however, the precision reduces as the number of prompts (test samples) increases. This is in direct contrast to common hypothesis testing, where larger sample size leads to smaller standard errors, thus higher precision. We get different results because our model provenance tester relies on detecting similarities of models. When using a smaller number of prompts, it can detect only the stronger similarities which are usually due to model provenance. However, as we increase the prompts, it starts detecting similar models that not necessarily have provenance relation. This leads to misclassification and reduced precision.
@@
We also examine the impact the randomness of prompt sampling on the tester’s accuracy. We conduct experiments on both benchmarks using five different randomly sampled sets of $1,000$ prompts [^3] We chose smaller number of prompts due to larger computation effort required to complete five full runs of both benchmarks. The running time is completely dominated by producing outputs from the models, which in theory is parallelizable, but in our case it was not due to limited GPU resources., with the same set of prompts used in all testers, and record the precision and recall for each run – see Table 6 of Appendix C. The results show that these values vary by $1-4\%$ between runs, indicating consistent performance across different prompt samples.
@@
The evaluations above are in the default framework, which assumes no candidate parent is given in each provenance test. We run similar experiments when the candidate parent is given in Appendix D. This is an easier problem (to make a wrong prediction, one needs not only to have conclusive hypothesis test that output a wrong parent, but also it should match the candidate parent), and the results confirm this: the recall of the tester in this framework is similar to the recall on the default framework, whereas the precision is very close to $100\%$.
@@ 4.2 Correctness of Assumptions
(RQ2): The assumption that derived models show significant similarity to their parent models is valid for most provenance pairs.
@@
Figure 3: Precision and recall of Bench-B (top) and Bench-A (bottom) with respect to smaller control set size.
(RQ3): The tester’s performance significantly degrades when the control set is too small or poorly selected.
@@ 4.3 Reducing Query Complexity
Certain pre-trained models from Bench-A and Bench-B exhibit a high degree of similarity when comparing their output tokens generated from random prompts. Table 7 of Appendix C presents the top $5$ most similar model pairs from Bench-B, measured by the percentage of matching output tokens when tested on 1,000 random prompts (column $k=1$).
@@
Figure 4: Recall for Bench-B with different values of advanced prompt sampling defined with $k$.
Figure 4 compares the tester’s recall across different values of $k$. Notable improvements are visible even at $k=4$, with higher values of $k$ showing better results (though with diminishing returns). Specifically, the recall achieved with $1,000$ prompts at $k=1$ can be matched using only about $250$ prompts at $k=64$, representing a four-fold reduction in online complexity. Figure 5 provides a comprehensive comparison between $k=1$ and $k=64$ for both precision and recall across both benchmarks, using $4-5$ times fewer queries for $k=64$ (note, in Figure 5 the number of prompts for $k=64$ are given at the top of the plots). The results demonstrate that the tester maintains its effectiveness despite the significant reduction in queries to the tested models. For example, advanced prompt sampling achieves high levels of $90-95$% precision and $80-90$% recall while reducing the required number of prompts from $3,000$ to just $500$ per model.
Figure 5: Comparison of precision/recall for Bench-B (top) and Bench-B (bottom) when advanced online prompt sampling with $k=64$ uses four times less prompts than no advanced sampling ($k=1$).
@@
5 Related Work
Model Ownership and Copyright Detection. The need for methods for determining model ownership and detecting illegitimate use of model is now well recognized; the main difference is in our formulation to tackle the concern. To determine model ownership, one type of techniques apply changes to the training dataset and model training in order to insert “watermarks” [24, 43, 53], backdoors [36, 2] or fingerprints [52, 8, 37, 28]. Specialized model ownership schemes have also been proposed for other models such as graph neural networks [55, 48]. All of these techniques, however, are orthogonal to the problem of provenance testing which we formulate in this paper. They require changes to the training of the parent model which may degrade the performance of the model. Moreover, most existing approaches do not have provable guarantees that model ownership can verified with a given confidence, so the verification is often empirically determined. Their focus is not on designing tests for determining provenance under model customizations. In particular, this is often a challenge for watermarking as they have limited robustness to typical model customizations, with some recent work acknowledging the challenge [26]. The closest work to ours is the recent work motivated by similar copyright licensing concerns [12]. The main difference to our approach is that we propose a tester with minimal assumptions of black-box access, while the prior work requires much more extra knowledge such as training and testing dataset and model parameters. Additionally, we evaluate and consider real-world customizations that are available on public repositories.
Hypothesis Tests in Machine Learning Security. Statistical hypothesis testing has been recently introduced for issues in machine learning security such as formalizing membership inference attacks via likelihood ratio tests [9, 41], auditing differentially private algorithms [30, 23], property inference [29], attribute inference [54, 17], or to proposing statistical verification methods for neural networks [6]. In this work, we also phrase the problem of provenance for fine-tuned LLMs as multiple statistical hypothesis testing, but our considered problem formulation is considerably different from those considered in prior works. There are several noteworthy differences in our formulation, compared to all of these other problem domains. In order to upper bound the hypothesis test’s false positive rate, we need to have a good estimate of the null hypothesis. For example, in some settings such as the membership inference, estimating the null hypothesis may be computationally expensive since it requires training different models under training datasets with or without given data points. Instead, in our formulation, we have multiple control models and we determine provenance depending on the reference similarity with them. As an example of another difference to privacy related inference tests, the randomness in our tests is only over the inputs given to the model, not training data points. As the choice of the input samples is independent from the training dataset of the model, our tests have soundness with respect to statistical significance.
Customization Techniques. Not many works study how fine-tuning or other customization techniques change the output distribution or the features of the pretrained model. Both training from scatch and fine-tuning optimize a similar training objective and they differ only in their initialization (random vs. pretrained weights). Because of the non-convex nature of the optimization, it is non-trivial to analytically analyze how their training dynamics converge to different minima. Some works theoretically analyze the effect of initialization (known as implicit regularization) for two-layer networks (or restricted setups), and not for pretraining [31, 42, 45]. Prior works denote that fine-tuning affects the robustness to out-of-distribution, pointing that some pretrained features get distorted in this process [27, 39, 44].

### v2

@@ 3.1 Model Provenance Tester
$\mathcal{M}\leftarrow\{\mu_{1},\ldots,\mu_{s}\}\cup\{\mu^{\prime}_{1},\ldots,\mu^{\prime}_{m}\}$ $\triangleright$ Set of all sims
$\mu_{max}\leftarrow\max(\mathcal{M})$ $\triangleright$ Find highest sim
@@
  $p_{i}\leftarrow\text{ZTest}(\mu_{max},\mu^{\prime},T)$ $\triangleright$ Compare against other sims
@@
for $k\leftarrow 1$ to s+m1s+m-1 do
  $\alpha_{k}\leftarrow\alpha/(s+m-k)$ $\triangleright$ Holm-Bonferroni adjustment
@@
end forreturn $(\textsc{True},\operatorname*{arg\,max}_{i\in[s]}\mu_{i})$ $\triangleright$ Return parent
The provenance tester described thus far assumes we have one candidate parent $f$ that we want to test our child model $g$ against. We now extend it to the general case where the goal is to determine whether a model $g$ is derived from some model from the set $f_{1},\ldots,f_{s}$} of candidate parent models. While running the basic tester $s$ times (once for each provenance pair $(f_{i},g)$) would solve the unspecified parent problem, it would also would require additional correction for multiple testing to maintain the same level of confidence. The probability of false positives would grow with the number of candidate parents $s$ unless appropriate adjustments are made to the significance level.
We thus propose the improved tester in Algorithm 1. Our tester avoids this issue by conducting a single set of hypothesis tests after identifying the most similar candidate. First, it finds the most similar model to the given model $g$ among all the control models $C$ and candidate parents $F$. If that model is a control model, the algorithm terminates with False. Otherwise, it goes on to test whether the FWER of this model is overall below $\alpha$, the desired significance level. When this test procedure returns False, it indicates that we could not establish this higher similarity with the desired level of statistical significance. This may occur either because there is genuinely no significant similarity indicative of provenance or because the test lacked sufficient power under the given parameters (e.g., sample size or number of prompts) to detect it. When the algorithm return True, it has the guarantee that the most similar model is one of the candidate models and that the total significance level across all hypotheses meets the threshold $\alpha$.
@@ 3.2 Understanding Sources of Error
The overall error rates of our tester depend on the combination of errors from both our statistical hypothesis testing and the two core assumptions which we describe below.
Assumption 1: Derivation implies similarity.  We assume that when model $g$ is derived from $f$, they will exhibit above-average similarity in their outputs. This assumption leads to two potential types of errors: false negatives and false positives. False negatives occur when a derived model shows insufficient similarity to its parent. This can happen when a model customizer applies extensive modifications that significantly alter the model’s behavior. While resource constraints typically prevent such extreme modifications (as they would approach the cost of training from scratch), some legitimate derivation relationships may still go undetected. False positives arise when independently developed models exhibit high similarity. This typically happens when models are trained on similar datasets or designed for similar specialized tasks - for instance, two independent medical diagnosis models may produce very similar outputs due to their shared domain constraints.
Assumption 2: Control models establish a valid baseline.  We assume our control models provide a reliable baseline for the similarity we should expect between unrelated models. Similarly, poor selection of control models can lead to two types of errors. False positives occur when our control models are too dissimilar from the domain of $f$. For example, using general language models as controls for specialized code generation models sets an artificially low baseline, making normal domain-specific similarities appear significant. False negatives happen when control models are themselves derived from $f$ or trained on very similar data. This establishes an artificially high baseline that masks genuine derivation relationships.
While we can provide theoretical guarantees for controlling error rates in hypothesis testing, we cannot derive analytical bounds for errors arising from the assumptions about derivation implying similarity or the validity of the control model baseline. The assumption-based error rates can only be evaluated empirically. We find these hold in practice through extensive experiments (Section 4).
@@ 3.3 Reducing Query Complexity
Most of LLMs available currently allow cheap (even free) API access, thus the monetary query cost of running our testers is insignificant. However, sometimes this is not the case, for instance when the cost of queries is high (e.g. one query to OpenAI model O1 can cost more than \$1 openai_pricing ), or when the models have some rate restrictions. Furthermore, there are use cases when query complexity can be reduced without any side effects, thus it makes sense from optimization perspective.
@@
We can divide the queries used in the tester into two distinctive types: online queries made to the tested child model $g$, and offline queries made to the parent model $f$ (or models $f_{1},\ldots,f_{s}$) and to the control models $c_{1},\ldots,c_{m}$. We make this distinction for two reasons. First, often offline queries are much cheaper, as the potential parent models (and the control models as we will see in the Section 4) are well established, and available from multiple sources, thus they are usually cheaper or free. Second, in some use cases, we can reuse the offline queries to perform many provenance tests of different $g_{i}$. We analyze separately these two scenarios.
@@
Reducing Online Complexity.  Since our tester is fundamentally based on statistical hypothesis testing, any reduction in query complexity must be compensated by increasing the statistical power of individual queries. Rather than querying model $g$ with $T$ random prompts, we can strategically select a smaller set of $T^{\prime}<T$ prompts that yield comparable statistical power for detecting model provenance [^1] It means in Algorithm 1, instead of random sampling $x_{1},\ldots,x_{T}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\Omega$, the goal is to find set $x_{1},\ldots,x_{T^{\prime}}$ from $x_{1},\ldots,x_{T}$ and $F,C$.. We achieve this through an informed sampling approach: instead of uniform sampling from $\Omega$, we employ rejection sampling with an entropy-based selection criterion. Specifically, to generate each prompt in $T^{\prime}$, we sample $k$ candidate prompts from $\Omega$ and select the one that maximizes the entropy of output tokens across all parent and control models. The selection criterion is dynamically weighted to favor prompts that have stronger discriminative power between similar models. While this approach introduces dependencies between the sampled prompts (so the theoretical guarantees of hypothesis testing used in Algorithm 1 do not carry over), our empirical results in Section 4.4 demonstrate its practical effectiveness. We detail our approach in Appendix H.
@@
Reducing Offline Complexity.  We further consider the case of reducing offline complexity in settings where offline queries cannot be reused. The key observation for reducing offline query complexity is that we may not need an equal number of queries to all parent/control models to identify the most similar one. If a particular parent model shows consistently higher similarity to $g$ compared to other models, we might be able to confirm it as the top candidate with fewer queries to the clearly dissimilar models. The challenge lies in determining when we have sufficient statistical evidence to conclude that one model is significantly more similar than the others, while maintaining our desired confidence levels. This observation naturally leads us to formulate the problem as a Best Arm Identification (BAI) audibert2010best problem in the Multi-Armed Bandit (MAB) setting. In this formulation, each parent or control model represents an “arm” of the bandit, and querying a model with a prompt corresponds to “pulling” that arm. The “reward” for each pull is the binary outcome indicating whether the model’s output matches that of the tested model $g$. The goal is to identify the arm (model) with the highest expected reward (similarity to $g$) while minimizing the total number of pulls (queries). We implement an optimization for our tester based on state-of-the-art BAI approaches (see Appendix I for details) and show it impact in Section 4.4.
@@ 4 Evaluation
We evaluate our proposed provenance testing approach experimentally and seek to answer the following research questions:
@@
4.1 Models and Provenance Testing Parameters
We collect model candidates for all provenance pairs from the Hugging Face (HF) platform huggingface . To avoid selection bias, we used download counts as our selection criterion, taking the most popular models subject only to hardware constraints on model size.
To increase variety of candidates, we create two distinct benchmarks Bench-A and Bench-B, that differ in aspects such as model sizes, choice of pre-trained models, and ground-truth verification procedure (refer to Tbl. 1). The full procedure of collection of models and constructions of benchmarks is described in Appendix C. We use the standard significance $\alpha=0.05$ (see Appendix G for other values). Sampling of prompts is given in Appendix D.
The implementation of the tester along with the two benchmarks can be found at https://github.com/ivicanikolicsg/model_provenance_testing.
@@ 4.1 Models and Provenance Testing Parameters
model parameters 1B-4B << 1B
compilation method manual automatic
@@
Selection of control set.  In all of our provenance tests, we use the complete set of available pre-trained models from the benchmark as control models - 10 models for Bench-A and 57 for Bench-B. This selection was done to demonstrates that effective control sets can be constructed without careful manual curation or domain-specific analysis. Manual curation would not have been feasible since our two benchmarks have over $600$ candidates. Specifically, we make no effort to align control models with particular parent models’ domains or capabilities. We neither analyze the outputs of parent models $f$ nor attempt to match control models to specific use cases. Instead, we simply include all pre-trained models that rank among the most popular on the Hugging Face platform. This sampling approach, while simple, helps avoid introducing obvious selection bias while ensuring broad coverage of model types and capabilities. Due to the large benchmark size, we find diverse examples of candidates for domains such as financial, medical and more (see Appendix C for details). We have not evaluated other selection strategies, which remains a direction for future work. This straightforward selection strategy, however, allows us to evaluate whether provenance testing can be effective even without carefully chosen control sets.
@@ 4.2 Accuracy of Model Provenance Tester
We evaluate the accuracy of the provenance tester by examining its performance on both Bench-A and Bench-B under different numbers of prompts. Figure 3 shows the precision and recall results from these experiments.
The precision is notably high (approximately $0.95$) when the tester uses up to $1,000$, and it is significantly higher than the baseline precision of randomly guessing the parent which is $\frac{1}{10+1}\approx 0.09$ for Bench-A and $\frac{1}{57+1}\approx 0.02$ for Bench-B. Interestingly, however, the precision reduces as the number of prompts (test samples) increases. This is in direct contrast to common hypothesis testing, where larger sample size leads to smaller standard errors, thus higher precision. We get different results because our model provenance tester relies on detecting similarities of models. When using a smaller number of prompts, it can detect only the stronger similarities which are usually due to model provenance. However, as we increase the prompts, it starts detecting similar models that not necessarily have provenance relation. This leads to misclassification and reduced precision.
@@
We also examine the impact the randomness of prompt sampling on the tester’s accuracy. We conduct experiments on both benchmarks using five different randomly sampled sets of $1,000$ prompts, with the same set of prompts used in all testers, and record the precision and recall for each run – see Table 7 of Appendix E. The results show that these values vary by $1-4\%$ between runs, indicating consistent performance across different prompt samples.
Figure 2: Precision and recall of Bench-B (left) and Bench-A (right) with respect to smaller control set size.
@@ 4.3 Correctness of Assumptions
(RQ2): The assumption of similarity between derived and parent models is valid for most provenance pairs.
@@
(RQ3): The tester’s performance degrades when the control set is too small or poorly selected.
@@ 4.4 Reducing Query Complexity
Figure 4: Recall for Bench-B with different values of advanced prompt sampling ($k$).
Figure 5: Precision/recall for Bench-B (left) and Bench-B (right) when advanced online prompt sampling with $k=64$ uses four times less prompts than no advanced sampling ($k=1$).
@@
Figure 4 compares the tester’s recall across different values of $k$. Notable improvements are visible even at $k=4$, with higher values of $k$ showing better results (though with diminishing returns). Specifically, the recall achieved with $1,000$ prompts at $k=1$ can be matched using only about $250$ prompts at $k=64$, representing a four-fold reduction in online complexity. Figure 5 provides a comprehensive comparison between $k=1$ and $k=64$ for both precision and recall across both benchmarks, using 454-5 times fewer queries for $k=64$ (note, in Figure 5 the number of prompts for $k=64$ are given at the top of the plots). The results demonstrate that the tester maintains its effectiveness despite the significant reduction in queries to the tested models. For example, advanced prompt sampling achieves high levels of 909590-95% precision and 809080-90% recall while reducing the required number of prompts from $3,000$ to just $500$ per model.